Find a power series for the function, centered at c, and determine the interval of convergence.(1+3x*) 2 h) f(x)= 3x4 – 5

Find a power series for the function, centered at c, and determine the interval of convergence.
Answer #1

(9) Given f(u) = (1+344) 2 ང Consider It 3x ५ AS power series of 20 (-1)?? Itx n=0 d 1 E co’lzy? It3 x 4 = no Differentiatiキー12x3 (-)”っ 410+1)-1 40+1) (サフェリー 00 9-125 (143898 {-13ht 302)(n+1) Rot3 か nao-) 1-3 *) <1 =) -123X=1 t = < 13 - 4 2 4 3 : Interval of converges to - I = 613) (-(i)t, (+) (-(ut () vots that -5/1-3xYn - - Z 2( X 3. 554 2 ~ 3n ㅋ 2 yn 3x43 - 2 63041) X Power Series of f(x)= 2 3xlg at C=o is f(x) = E 2. 263 yo x got 3 ffr) s3. < x < - 3/3 < x < 3 -34 ( ut jos 176-1931) u Interval of convergence = (ty to